solve each system by elimination. check your answers.

-x-y-2z=4
-x+2y+z=1
-x+y-3z=11

Answer by checkley75(3666)

-x-y-2z=4
-x+y-3z=11 now add.
--------------------------------
-2x-5z=15
now multiply equation 1 by 2
-2x-2y-4z=8
-x+2y+z=1 now add
--------------------------------------
-3x-3z=9 now multiply this equation by 2
-6x-6z=18 now multiply -2x-5z=15 by -3 & add
6x+15z-45
--------------------
9z=-27
z=-27/9
z=-3 answer.
-3x-3*-3=9
-3x+9=9
-3x=0
x=0 answer.
0-y-2(-3)=8
0-y+6=8
-y=8-6
-y=2
y=-2 answer.